OpenStudy (anonymous):

the probability of getting a white marble is 1/5. if there are 37 marbles in the bag how many total marbles is there?

jimthompson5910 (jim_thompson5910):

P(White) = (number of white)/(number total) 1/5 = 37/x solve for x

OpenStudy (anonymous):

I need a explanation though, its what i dont understand.

jimthompson5910 (jim_thompson5910):

cross multiply to get 1*x = 5*37 x = ???

jimthompson5910 (jim_thompson5910):

probability is just a fraction of (number you want)/(number total) you would set the probability fraction equal to the given fraction 1/5 and solve for x